33, 48

Depicted below are a peptide substrate and the three catalytic residues within the chymotrypsin active site Note that the 2-dimensional geometry is only roughly shown. R1 represents a large hydrophobic amino acid and R represents other amino acids. The R groups of the catalytic triad of aspartate (Asp102), serine (Ser195), and histidine (His57) are shown. Select the atoms that act as nucleophiles and electrophiles in the acylation phase of the reaction Scroll down for more of this question Note: Selected atoms will turn green (undo by clicking again) Select the 2 atoms that act as nucleophiles in the acylation phase. Note: As shown below, the nucleophilic atoms might be weak nucleophiles, but they become strong in the process of catalysis Select the 2 atoms that act as electrophiles in the acylation phase NH NH Peptide substrate Asp102 His57 Ser195 Asp102 His57 Ser195 Peptide substrate
Answer #1

- OH group of serine acts as nucleophile. C=O of peptide substrate acts as electrophile in the catalysis.

Molarity is given by

Number of moles in per Litre of solution or in other words

Mass /molecular weight *1/volume in Litre.

0.550 *10-6 =23*10-3/molecular weight *1000/10

Molecular weight =23*10-3*1000/(5.5 *10-7*10)

23*106/5.5

4.16*106

4160kDa

Given that Ag reacts with sulfhydrly group. So assuming that there is only one sulfhydryl group is present in the protein then only we can say that the molecule weight is the minimum molecular weight. If it had more that one then molecular weight will be the multiple of what we have calculated.

So the answer is option a